OpenStudy (anonymous):

Solve the ODE x(dy/dx)=y-(x^2+y^2)^1/2

sam (.sam.):

dy/dx = (y/x) - √(1 + (y/x)²). So set y = xu. Then dy/dx = x du/dx + u ==> x du/dx + u = u - √(1 + u²) ==> du/√(1 + u²) = - dx/x. You can use a trigonometric substitution to integrate the left side. Putting u = tanΘ gives ∫ du/√(1 + u²) = ∫ secΘ dΘ = ln|secΘ + tanΘ| = ln|√(1 + u²) + u|. ln|√(1 + u²) + u| = - ln|x| + C ==> ln|√(1 + u²) + u| + ln|x| = C Use properties of logs (ln(a) + ln(b) = ln(ab)) and substitute u = y/x to get ln|√(x² + y²) + y| = C ==> √(x² + y²) + y = A where A = e^C.

OpenStudy (anonymous):

Hi, I need help on this questions also which I posted as doc file.

OpenStudy (anonymous):

Thanks a lot.. :)